### v1

# Macroscopic States in Bose–Einstein Condensate Dark Matter Model with Axionlike Interaction

Abstract
The phase diagrams of ultralight dark matter (DM), modeled as a self-gravitating Bose–Einstein condensate with axionlike interaction, are studied. We classify stable, metastable and unstable DM states over a wide range of condensate wave function amplitudes. It is shown that the axionlike interaction causes instability and an imaginary speed of sound at low amplitudes, whereas in a specific high-amplitude band, DM attains a stable state capable of forming a dense solitonic core and suppressing quantum fluctuations in the surrounding galactic DM halo. These findings are corroborated by evaluating thermodynamic functions for DM in the dwarf galaxy NGC 2366 and in its hypothetical analogues with different core-to-halo mass ratios. Distinct DM phase compositions respond differently to fluctuation-induced partial pressure, resulting in a first-order phase transition in a certain range of an interaction parameter. While the DM properties in NGC 2366 lie within the supercritical regime, the phase transition nonetheless provides a thermodynamic marker separating stable from unstable DM configurations. Once a dense core forms–reaching a threshold of about 12% of the total mass–the enhanced gravitation stabilizes the DM halo against fluctuations, while the internal pressure ensures core stability. In particular, we find that NGC 2366’s dense DM comprises roughly 19% of the DM mass while occupying only 4.7% of its total volume.
axionlike dark matter, dwarf galaxies, Bose–Einstein condensate, phase transitions
pacs:
95.35.+d, 05.30.Jp, 05.70.Ce, 05.70.Fh
I Introduction
Several decades ago, axions were proposed as hypothetical pseudo-Goldstone bosons PQ77 , offering a solution to the charge-parity (CP) problem in quantum chromodynamics (QCD). This innovative idea required an explanation of both the emergence of axions during the Universe’s evolution and their broader cosmological implications. Axions were soon regarded as promising candidates for dark matter (DM) PWW . The physical picture was theoretically supported by mechanisms for their production AS83 ; DF83 ; Dav86 and their transition to the nonrelativistic condensate state SY09 ; SY12 ; Dav13 ; Dav15 ; GHP15 . The latter allowed to describe low-energy axions with large occupation numbers either as a classical field in effective field theory or as a macroscopic wave function in the Gross–Pitaevskii approach.
Often considered within the context of DM structures, modeling of axion stars and clumps has revealed that their masses and sizes are relatively small from a cosmological perspective (see BZ19 ) and are constrained by the gravitational stability conditions inherent to cold bosonic matter BGZ84 ; UTB02 . As argued in Bor16 ; KM17 ; FMT08 ; Chavanis2 , the masses $10^{-4\pm 1}~{}\text{eV}c^{-2}$ of axions preclude forming giant condensates comparable in size to galactic DM halos.
In contrast, early attempts to model DM halos and rotation curves of galaxies yielded extremely small values of the boson masses Bal83 ; Sin that are tens of orders of magnitude smaller than those predicted in QCD. Therefore, reproducing the observed data requires invoking dark ultralight axions or axionlike particles Mar16 .
Although the presence of self-interactions between ultralight DM particles ensures a better description of astrophysical phenomena, the fuzzy DM models often omit it (see Hu ; Schive1 ; Chen17 ; Zim25 ), limiting themselves to taking into account only gravitational interactions, as was originally assumed Zwicky .
It seems natural for us to investigate self-gravitating DM in real visible objects using dark bosons with masses $10^{-22\pm 1}~{}\text{eV}c^{-2}$, incorporating the simplest instanton self-interaction borrowed from QCD BZ19 . Specifically, our goal is to apply the equilibrium statistical approach to the DM in dwarf galaxies, effectively replacing, in an ergodic sense, the analysis of the time evolution of a single spatial configuration of DM by studying a statistical ensemble of static configurations, which is governed by thermodynamic potential. In this way we intend to classify and describe the phases of such a matter.
Axionlike interactions can give rise to both stable and unstable states, allowing their coexistence in nonuniform DM. It is already known that the unstable state arises in the self-attraction mode Nambo24 at small amplitudes of the axion field, corresponding to rarefied matter. Although high amplitudes seem unlikely for an axion field (but not necessarily for other substances), some of their values should correspond to stable states KT94 . In any case, the overall state of the DM is determined by the spatial distribution and relative abundance of these states, which collectively define DM composition. Accordingly, the complete configurations are also classified as either stable or unstable, and it is assumed that these regimes can change via a first-order quantum phase transition at zero temperature.
One may observe parallels with the ensemble of oscillons Osc1 generated by the instanton self-interaction, leading in the absence of gravity to the three-dimensional sine-Gordon equation Man04 . Moreover, established results on axitons BZ19 ; KT94 ; ax5 are valuable for further comparison of sets of states.
Neglecting the Universe’s evolution in describing current visible structures, we thus focus on the statistical properties of self-interacting DM (SIDM) within the model of self-gravitating Bose–Einstein condensate (BEC). Phase diagrams should then reflect these properties, tracing transitions among distinct states and configurations.
Numerous studies have rigorously examined the properties of BEC DM with two- and three-particle interactions, using both exact numerical methods and various analytical approximations, such as the Thomas–Fermi approximation and the variational method (see Chavanis2 ; Peebles ; Bohmer ; Harko2011 ; Zang2018 ; Nazar ; SV24 ). The thermodynamic functions in these models have been carefully studied, and a new aspect of DM has emerged GKN20 ; GN21 ; GN23 : the prediction of a first-order phase transition between rarefied and dense states, which is driven by quantum fluctuation pressure and governed by coupling constants. Additionally, we would like to note the impact of the formation of composites from the initial DM particles GN23 ; GN22 ; GKKN , as well as the quantum entanglement of parts of the DM halo GN21 ; Lee18 .
Building on those studies, we transition from polynomial-type self-interaction (over a field or wave function) to a nonlinear periodic one.
To constrain the parameters of our BEC DM model described by the Gross–Pitaevskii–Poisson equations, we consider the DM dominating dwarf galaxy NGC 2366 as a reference. This nearby galaxy provides an opportunity to study reionization processes in vigorous low-metallicity starbursts, such as those that occurred during the cosmic reionization epoch Iz99 . In this regard, we intend to calibrate the model by fitting its observed rotation curve Oh08 , and then generate a statistical ensemble of hypothetical DM distributions following from the model equations under similar conditions. By comparing different configurations using phase diagrams, we aim to reveal the uniqueness of the DM structure in NGC 2366, which forms the gravitational background for subgalactic processes.
The paper is structured as follows. In Sec. II we introduce the main equations of the model and, through various simplifications, analyze the properties essential for interpreting subsequent findings. In Sec. III we explore nonperturbative solutions and their application to modeling the DM distribution in the dwarf galaxy NGC 2366. We then present our main results – statistical characteristics and phase diagrams based on the numerically obtained DM profiles. The final Sec. IV contains a summary of the findings and concluding remarks.
II The model
II.1 The model equations
Consider a stationary Bose–Einstein condensate (BEC) of nonrelativistic particles with mass $m$, described by a macroscopic real wave function $\psi(r)$ of a radial coordinate $r$ within a three-dimensional ball $B(r\leq R)$. In the absence of hydrodynamic flows, the incorporation of the axionlike self-interaction $V_{\rm int}$ and the potential of gravitational interaction $V_{\rm gr}$ leads to the energy functional $\Gamma=E-{\tilde{\mu}}{\cal N}$ with a constant chemical potential $\tilde{\mu}$:
$\Gamma=4\pi\int_{0}^{R}\widetilde{\cal F}r^{2}\,\mathrm{d}r,$ (1)
~$\widetilde{\cal F}=\frac{\hbar^{2}}{2m}(\partial_{r}\psi)^{2}+mV_{\rm gr}\psi^{2}+V_{\rm int}-\tilde{\mu}\psi^{2},$ (2)
$V_{\rm int}=\frac{U}{v}\left[1-\cos{\left(\sqrt{v}\psi\right)}\right]-\frac{U}{2}\psi^{2},$ (3)
$\Delta_{r}V_{\rm gr}=4\pi Gm\psi^{2}.$ (4)
The spatial evolution of $\psi(r)$ is governed by the minimization of the functional $\Gamma$ and the Poisson equation (4). This equation, along with its solution, is based on the radial part of the Laplace operator $\Delta_{r}$ and its inverse $\Delta^{-1}_{r}$ with respect to the variable $r$:
$\Delta_{r}f(r)=\partial^{2}_{r}f(r)+\frac{2}{r}\,\partial_{r}f(r),$ (5)
$\Delta^{-1}_{r}f(r)=-\frac{1}{r}\int_{0}^{r}f(s)s^{2}\mathrm{d}s-\int_{r}^{R}f(s)s\mathrm{d}s.$ (6)
The axionlike interaction used here is chosen in analogy with PQ77 ; Wit80 , assuming the connection between the axion field $\varphi$ with the nonrelativistic wave function $\psi$ via the relation $|\varphi|^{2}=(\hbar^{2}/m)|\psi|^{2}$ Chavanis2 . Other effective axion potentials also appear in VV80 ; Guzman ; Cor16 ; Chavanis3 .
Expanding (3) in a series, we observe that the first term $\psi^{4}$ corresponds to a two-particle self-interaction, while the subsequent term $\psi^{6}$ indicates a three-particle self-interaction. The BEC DM models with only pairwise repulsive interaction at $v=-|v|$ have been intensively studied in the Thomas–Fermi approximation (e.g. see Harko2011 ; Zang2018 ; Nazar ). Considering the three-particle interaction alongside quantum fluctuations substantially enriches the physical picture, leading to a two-phase structure and a first-order phase transition GKN20 ; GN21 .
The axionlike interaction (3) is characterized by two constants, $U$ and $v$, which have dimensions of energy and volume. In the particle physics VV80 ; Cor16 , they are related with the axion mass $m$ and decay constant $f_{\text{a}}$ as $U=mc^{2}$, $v=\hbar^{3}c/(mf^{2}_{\text{a}})$. Although the applicability of these relativistic relations in nonrelativistic models is controversial, they allow us to make comparisons.
In cosmological models, axions are described through various parameterizations Chavanis2 . As highlighted in FMT08 , the choice of a smaller particle mass ensures the formation of certain structures in the Universe, some of which we aim to describe. It is expected that $m\sim 10^{-22}~{}\text{eV}c^{-2}$, which is usually attributed to the fuzzy DM Hu ; Peebles .
Then, introducing the distance scale $r_{0}$, in nonrelativistic quantum picture we fix $U$ as
$U=\frac{\hbar^{2}}{mr_{0}^{2}}.$ (7)
Although the relativistic value $U=mc^{2}$ is recovered at the Compton wavelength $r_{0}=\hbar/(mc)$, the chosen $U$ allows equalizing the scales of quantum fluctuations and self-interaction, similar to the sine-Gordon model Man04 ; Wil18 . In any case, the implications of changing $U$ are discussed below.
It immediately leads to the dimensionless measure of gravitational interaction:
$A=\frac{8\pi Gm^{3}r_{0}^{4}}{\hbar^{2}v}\ll 1.$ (8)
With $v=\hbar^{3}c/(mf^{2}_{\text{a}})$ defined via the axion decay constant $f_{\mathrm{a}}$, it is expected that in models featuring ultralight particles, $f_{\text{a}}$ may fall within the range $10^{18}\,\text{eV}<f_{\text{a}}<10^{21}\,\text{eV}$ Chavanis2 .
Defining the dimensionless radial variable $\xi$ and wave function $\chi$,
$\xi=\frac{r}{r_{0}},\hskip 25.0pt\chi(\xi)=\sqrt{v}\,\psi(r),$ (9)
we reduce the problem (1)-(4) to the following:
$\Gamma=\Gamma_{0}\int_{0}^{\xi_{B}}{\cal F}\xi^{2}\,\mathrm{d}\xi,\quad\Gamma_{0}=\frac{2\pi r_{0}f^{2}_{\text{a}}}{\hbar c},$ (10)
${\cal F}=(\partial_{\xi}\chi)^{2}+A\Phi\chi^{2}+2(1-\cos{\chi})-\nu\chi^{2},$ (11)
$\Delta_{\xi}\Phi=\chi^{2},\quad\Phi(0)=0.$ (12)
It operates on a set $r_{0}$, $f_{\text{a}}$, $A$ and $\nu$, such that
$m=\left[\frac{\hbar^{5}c}{8\pi G}\frac{A}{r_{0}^{4}f^{2}_{\text{a}}}\right]^{1/4},\quad\rho_{0}=\frac{m}{v},$ (13)
$\nu=1+2u-A\Phi_{0},\hskip 22.76219ptu=\frac{mr_{0}^{2}}{\hbar^{2}}\tilde{\mu},$ (14)
where $\rho_{0}$ is the mass density scale, and $\nu$ plays the role of the effective chemical potential, which absorbs part of the axion interaction and the gravitational potential at the origin $\xi=0$, namely
$\Phi_{0}=-\int_{0}^{\xi_{B}}\chi^{2}(\xi)\,\xi\,\mathrm{d}\xi.$ (15)
In our study, $\nu$ is treated as a free parameter, reflecting the arbitrariness of $\tilde{\mu}$.
Thus, the DM characteristics are
$m$ $=$ $0.17745\times 10^{-22}~{}\text{eV}~{}c^{-2}~{}\left[\frac{A}{10^{-3}}\right]^{1/4}$ (16)
×$\times\left[\frac{r_{0}}{1~{}\text{kpc}}\right]^{-1}\left[\frac{f_{\text{a}}}{10^{19}~{}\text{eV}}\right]^{-1/2},$
$\rho_{0}$ $=$ $0.73055\times 10^{-23}~{}\text{kg}~{}\text{m}^{-3}~{}\left[\frac{A}{10^{-3}}\right]^{1/2}$ (17)
×$\times\left[\frac{r_{0}}{1~{}\text{kpc}}\right]^{-2}\left[\frac{f_{\text{a}}}{10^{19}~{}\text{eV}}\right],$
so that the central mass density $\rho_{c}=\rho_{0}\chi^{2}(0)$.
The energy scale is defined as $\varepsilon_{0}=\hbar^{2}/(2mr_{0}^{2})$. Combining this with the mass scale, we get
$\varepsilon_{0}$ $=$ $1.1522\times 10^{-30}~{}\text{eV}~{}\left[\frac{A}{10^{-3}}\right]^{-1/4}$ (18)
×$\times\left[\frac{r_{0}}{1~{}\text{kpc}}\right]^{-1}\left[\frac{f_{\text{a}}}{10^{19}~{}\text{eV}}\right]^{1/2}.$
After the manipulations carried out, we come to the model equations in terms of the wave-function $\chi(\xi)$ and gravitational potential $\Phi(\xi)$:
$\left(\Delta_{\xi}+\nu\right)\chi-A\Phi\chi-\sin{\chi}=0,$ (19)
$\Phi(\xi)=-\frac{1}{\xi}\int_{0}^{\xi}\chi^{2}(s)s^{2}\,\mathrm{d}s+\int_{0}^{\xi}\chi^{2}(s)s\,\mathrm{d}s,$ (20)
where $\Delta_{\xi}\Phi=\chi^{2}$ is guaranteed, and $\Phi(0)=0$.
Note the scaling symmetry of Eqs. (19)-(20). Let $\widetilde{\chi}(\xi)$ be a solution to the set of equations:
$\left(\Delta_{\xi}+\widetilde{\nu}\right)\widetilde{\chi}-\widetilde{A}\Phi\widetilde{\chi}-\widetilde{B}\sin{\widetilde{\chi}}=0,$ (21)
$\Phi(\xi)=-\frac{1}{\xi}\int_{0}^{\xi}\widetilde{\chi}^{2}(s)s^{2}\,\mathrm{d}s+\int_{0}^{\xi}\widetilde{\chi}^{2}(s)s\,\mathrm{d}s,$ (22)
Suppose there exists a constant $\alpha$ such that $\widetilde{\nu}=\nu\alpha^{2}$, $\widetilde{B}=\alpha^{2}$ and $\widetilde{A}=A\alpha^{4}$. Then, provided that the initial conditions agree $\widetilde{\chi}(0)=\chi(0)$, the solutions to Eqs. (19)-(20) and (21)-(22) are related by $\widetilde{\chi}(\alpha\xi)=\chi(\xi)$. This means that replacing any $\widetilde{B}>0$ by one changes the scale of $\xi$ and hence $r_{0}$, which partially justifies using only the parameters $\nu$ and $A$ as independent.
In what follows we will focus on the analysis and solution of Eqs. (19)-(20), where the first of them serves as the stationary Gross–Pitaevskii equation.
II.2 The particular cases
Uniform states. Let us examine the simplest scenario – a nonfluctuating BEC DM without self-gravity. In this case, when $\partial_{\xi}\chi=0$ and $A=0$, the grand canonical potential density (i.e. per unit volume) is
$\omega=2(1-\cos{\chi})-\nu\chi^{2},$ (23)
where $\nu$ is a chemical potential or external field.
The particle number density $\eta$ and volume per particle $v$ are
$\eta\equiv-\frac{\partial\omega}{\partial\nu}=\chi^{2},\quad v=\frac{1}{\eta}.$ (24)
Figure 1: Grand canonical potential density $\omega$ as a function of order parameter $\chi$ for various $\nu$. The curve numbers $i=\overline{1,5}$ correspond to $\nu_{i}=b_{i}$, see (32), and curve 6 is for $\nu=0$. The two dotted curves passing through the minima of $\omega$ define the ranges of existence of phase-1 and phase-2.
Within the range of variation of the order parameter $\chi$ depicted in Fig. 1, all curves $\omega(\chi)$ exhibit a minimum at $\chi=0$. However, for $\nu>1$, the value $\chi=0$ determines unphysical maximum, making curve 1 a boundary curve. At $\nu=0$, curve 6 possesses infinitely many identical minima of $\omega$ at $\chi=2\pi n$ (where $n\in\mathbb{N}$), corresponding to coexisting uniform phases.
We limit ourselves by considering the two distinct phases in curve 6: the phase-1 at $\chi=2\pi$ and the denser phase-2 at $\chi=4\pi$. These pure phases manifest within specific intervals of variation for $\nu$ and $\chi$. We have found that the phase-1 exists for $\nu\in[\nu_{2};\nu_{3}]$, while the phase-2 is present within the range $\nu\in[\nu_{4};\nu_{5}]$.
The parameters $\nu_{1}$-$\nu_{5}$ are (locally) critical for the curves in Fig. 1. Their values are established through conditions:
$\frac{\partial\omega}{\partial\chi}=0,\hskip 25.0pt\frac{\partial^{2}\omega}{\partial\chi^{2}}=0.$ (25)
For further analysis, we utilize the spherical Bessel functions:
$j_{l}(z)=z^{l}\left(-\frac{1}{z}\frac{\mathrm{d}}{\mathrm{d}z}\right)^{l}\frac{\sin{z}}{z},$ (26)
so that $j_{l}(0)=0$ for $l\geq 1$.
We also need the extrema of $j_{0}(z)$. Since $j_{0}^{\prime}(z)=-j_{1}(z)$, let us define the auxiliary sets of numbers $\{a_{n}\}$ and $\{b_{n}\}$, where $n\in\mathbb{N}$, by $j_{1}(a_{n})=0$ and $b_{n}\equiv j_{0}(a_{n})$. The first five elements of the sequences are
$\begin{array}[]{ll}a_{1}=0,&\ b_{1}=1;\\ a_{2}=4.493409,&\ b_{2}=-0.2172336;\\ a_{3}=7.7252518,&\ b_{3}=0.12837455;\\ a_{4}=10.90412,&\ b_{4}=-0.0913252;\\ a_{5}=14.06619,&\ b_{5}=0.070913459.\end{array}$ (32)
Thus, rewriting (25) as
$\nu=j_{0}(\chi),\hskip 25.0ptj_{1}(\chi)=0,$ (33)
one finds the set of solutions: $\chi_{i}=a_{i}$ and $\nu_{i}=b_{i}$.
Note that the equation $\nu=j_{0}(\chi)$ with respect to $\chi$ can have several solutions, as shown in Fig. 2.
Figure 2: Solution to the equation $j_{0}(\pi x)=j_{0}(\pi y)$ in the domain $x,y\in[0;5]$. Red sections correspond to positive values, $j_{0}(\pi x)\geq 0$, while the blue ones are for $j_{0}(\pi x)<0$.
After determining the ranges in which pure phases exist, we examine their manifestations by turning to thermodynamic functions. The energy density $e$ and free energy density $f$ coincide at zero temperature to be
$f=e=\omega+\nu\eta=2(1-\cos{\chi}).$ (34)
Then the chemical potential $\mu$ and the internal pressure $p$ are
$\mu$$\equiv$ $\frac{\partial f}{\partial\eta}=j_{0}(\chi),$ (35)
$p$$\equiv$ $-\frac{\partial(vf)}{\partial v}=\eta\frac{\partial f}{\partial\eta}-f$ (36)
$=$ $\chi\sin{\chi}-2(1-\cos{\chi}).$
Therefore, the equation $\partial\omega/\partial\chi=0$ implies the condition $\mu(\chi)=\nu$ as before. Then all extrema of the potential $\omega$ in Fig. 1 are linked by the single function $\widetilde{\omega}(\chi)=-p(\chi)$, so that the pink and blue dotted curves are its fragments.
Note that the chemical potential $\mu$, internal pressure $p$ and internal energy density $e$ obey the Duhem–Gibbs and Euler relations at zero temperature LL :
$\mathrm{d}p=\eta\,\mathrm{d}\mu,\hskip 25.0pte=\mu\eta-p.$ (37)
Expanding $p$ for relatively small density $\eta$, the equation of state takes on the form:
$p=-\frac{\eta^{2}}{12}+\frac{\eta^{3}}{180}+O(\eta^{4}).$ (38)
The terms $-\eta^{2}$ and $\eta^{3}$ can be related to two-particle attraction and three-particle repulsion. The interplay of these effects may induce a first-order phase transition, although such a transition is anticipated at higher densities.
Figure 3: Chemical potential $\mu$ versus internal pressure $p$, which are parametrized by $\chi$ in the range of $\chi\in[0;4.5\pi]$. Segments $AB$ and $CD$ are unstable states; $BO$ and $OE$ represent stable states of phase-1 and phase-2; $OC$ and $DO$ correspond to metastable states of phase-1 and -2. The first-order phase transition between phase-1 and -2 occurs at point $O$.
Let us analyze the $p-\mu$ phase diagram in Fig. 3, which reflects all states for $\chi\leq 4.5\pi$. Guided by the minimum of potential $\omega$, as shown in Fig. 1, we conclude that for $\mu<0$, the phase-1 is favored over the metastable denser phase-2, whereas for $\mu>0$, the phase-2 becomes dominant. At the same time, Fig. 3 also shows unstable states $AB$ and $CD$ corresponding to local maxima of the potential $\omega$, when
$\frac{\partial\omega}{\partial\chi}=0,\hskip 25.0pt\frac{\partial^{2}\omega}{\partial\chi^{2}}<0.$ (39)
Considering the parametric $p-\mu$ dependence for arbitrarily large $\chi$, one readily finds that the value $\chi_{c}\equiv a_{2}$ indicates the threshold for the emergence of stable phase in BEC DM. We stress once again that the interval $\chi\in[0;\chi_{c}]$ – depicted as segment $AB$ in Fig. 3 – represents unstable state.
We stress that the simplifications adopted in the small-$\chi$ regime restrict the analysis to the unstable state only, thereby missing the complex – physically rich – picture of phenomena.
In Fig. 3, the turning points at $\chi=a_{n}$, where the phase pairs converge, exhibit divergent compressibility, as determined by the expression:
$-\frac{\mathrm{d}\chi}{\mathrm{d}p}=\frac{1}{\chi^{2}j_{1}(\chi)}.$ (40)
This property is sometimes used to identify quantum phase transitions NT16 . It is also related to the (dimensionless) speed of sound, the square of which is
$c^{2}_{s}(\chi)=\frac{\mathrm{d}p}{\mathrm{d}\eta}=-\frac{\chi}{2}j_{1}(\chi).$ (41)
We see that $c^{2}_{s}(a_{n})=0$ and $c^{2}_{s}(\chi)<0$ in unstable states, for example, for $\chi\in[0;\chi_{c}]$. Note that similar cases are known in the laboratory neg.sss-1 ; neg.sss-2 ; neg.sss-3 .
If the admissible range of the order parameter $\chi$ does not exceed $\chi=a_{3}$, implying that no phases other than phase-1 exist, then the entire segment $BC$ of phase-1 can be regarded as stable.
Without loss of generality, let us shift the potential $\omega$ by an arbitrary constant, thereby redefining its origin:
$\omega=-p_{0}-2\cos{\chi}-\nu\chi^{2}.$ (42)
This leads to the corrected functions:
$p=p_{0}+\chi\sin{\chi}+2\cos{\chi},$ (43)
$e=-p_{0}-2\cos{\chi},$ (44)
while maintaining the form of $\mu=j_{0}(\chi)$ and (37).
The value of constant $p_{0}$ for an inhomogeneous body in a vacuum should be determined from the conditions $p(\chi_{B})=0$ and $\mu(\chi_{B})=0$ for a boundary value $\chi_{B}$. It follows from them that $\chi_{B}=\pi k$ and $p_{0}=-2\cos{(\pi k)}$ for $k\in\mathbb{Z}$. Thus, $p_{0}=-2$ for $k=2n$ or $p_{0}=2$ for $k=2n+1$, $n\in\mathbb{Z}$.
Note that the value $p_{0}=-2$ ensures the convergence of integrals in the spatially one-dimensional sine-Gordon theory Man04 . In contrast, we consider a three-dimensional model in a finite spherical volume.
Thomas–Fermi approximation. It involves disregarding the contribution of quantum fluctuations $\Delta_{\xi}\chi$ in Eq. (19), while considering the gravitational interaction, which leads to the set:
$j_{0}(\chi)=\nu+\frac{A}{\xi}\int_{0}^{\xi}\upsilon(s)\,\mathrm{d}s,\quad j_{0}(\chi(0))=\nu,$ (45)
[…]
where the gravitational potential $\Phi(\xi)$ is identically re-written as
$\Phi(\xi)=\frac{1}{\xi}\int_{0}^{\xi}\upsilon(s)\,\mathrm{d}s,\quad\upsilon(\xi)=\int_{0}^{\xi}\chi^{2}(s)\,s\,\mathrm{d}s.$ (47)
Figure 4: Wave functions at $A=4.5\times 10^{-3}$ and $\chi(0)=6.551147852$. Red line is the solution of Eqs. (45)-(46) at $\nu=0.04041539748$, and it terminates at $\xi_{B}=3.0747465$, indicated by the vertical dashed line. Blue curve represents the solution of Eqs. (19)-(20) at $\nu=0.36$.
Typical numerical solution of Eqs. (45)-(46) is shown in Fig. 4 (red line). It describes a drop of BEC DM in the phase-1. Physically, it is assumed that at the origin, the dense phase is concentrated, exhibiting a predominance of self-repulsion. Therefore, we take $2\pi<\chi(0)<2.5\pi$, according to Fig. 3, to get $j_{0}(\chi(0))>0$.
To find out the reason for the discontinuity of the solution, let us substitute the gravitational potential to obtain a closed equation for the wave function $\chi$:
$j_{1}(\chi)\Delta_{\xi}\chi+\left(\chi^{\prime}\right)^{2}\left[\frac{j_{1}(\chi)}{\chi}-j_{2}(\chi)\right]+A\chi^{2}=0.$ (48)
where $\chi^{\prime}(\xi)\equiv\partial_{\xi}\chi(\xi)$.
It is easy to see that the spatial development stops when the first term vanishes due to $j_{1}(\chi)=0$, which happens at $\chi=a_{2}$. Such $\chi$ is actually reached at $\xi_{B}$ in Fig. 4. At $\xi>\xi_{B}$ we have to put $\chi=0$, but the gravitational potential behaves as $-1/\xi$.
In principle, integration would be interrupted for any model at $\chi(\xi_{B})=a_{i}$ if $\chi(0)\in(a_{i};a_{i+1})$, where the five numbers of $\{a_{n}\}$ are shown in (32).
Since the nonrelativistic gravitational potential breaks the symmetry with respect to the shift $\chi\mapsto\chi+2\pi$, let us analyze the side effects. Note that for $V_{\rm gr}=0$ the invariance of the functional (1) in terms of dimensionless variables is ensured, in particular, by the given pseudoscalarity of $\chi$.
So, let us write the wave function as $\chi=2\pi+\widetilde{\chi}$ and consider the Poisson equation:
$\Delta_{\xi}\Phi=\chi^{2},\hskip 25.0pt\Phi(0)=0.$ (49)
We immediately see that the extracted constant contribution $4\pi^{2}$ to density produces such a term in $\Phi$:
$\Phi_{\rm rot}(\xi)=\frac{2\pi^{2}}{3}\xi^{2}.$ (50)
Its contribution to Eq. (19) may be associated with a rotation, the role of which is already noted (see Zang2018 ; Nazar ). If the system remains unbounded, nonrelativistic rotation with increasing $\xi$ produces a divergent, node-free solution $\chi(\xi)$, signaling inherent instability. However, periodic self-interaction of limited magnitude may be insufficient to arrest this divergence. The finite part of such a solution to (19)-(20), the search of which is discussed below, is plotted in Fig. 4 (blue curve). Nevertheless, similar solutions turn out to be suitable for describing a number of observed data (see Fig. 6).
Moreover, we refer to the model from GKN20 , which suggests to separate the approaches to describing the core and the tail of the DM halo due to different role of self-interaction in relatively dense and rarefied regions.
II.3 The excitation spectrum and instabilities
Given the quantum-fluid nature of BEC DM, let us analyze the spectrum of its small excitations. We reformulate Eq. (19) into the non-stationary Gross–Pitaevskii equation for the complex-valued wave function $X$, which depends on time $t$, as
$\left(-\Delta_{\xi}+W\right)X=\mathrm{i}\partial_{t}X,$ (51)
$W(\xi,t)=A{\cal V}(\xi,t)+j_{0}(|X(\xi,t)|),$ (52)
${\cal V}(\xi,t)=-\frac{1}{\xi}\int_{0}^{\xi}|X(s,t)|^{2}s^{2}\,\mathrm{d}s$
$\hskip 39.83385pt-\int_{\xi}^{\xi_{B}}|X(s,t)|^{2}s\,\mathrm{d}s,$ (53)
where $\mathrm{i}=\sqrt{-1}$.
To derive both the continuity equation and the Euler hydrodynamic equation, we substitute $X=\chi\exp{(\mathrm{i}\theta)}$, where $\chi$ and $\theta$ are real functions of space and time.
Substituting the wave function $X$ into (51) and separating the imaginary and real parts, we obtain the desirable equations:
$\partial_{t}\eta+{\boldsymbol{\nabla}}(\eta{\bf v})=0,$ (54)
$\partial_{t}{\bf v}+{\boldsymbol{\nabla}}\left(\frac{{\bf v}^{2}}{2}+2W-\frac{2}{\sqrt{\eta}}\Delta_{\xi}\sqrt{\eta}\right)=0,$ (55)
where the velocity vector ${\bf v}\equiv 2{\boldsymbol{\nabla}}\theta$ is expressed via the gradient operator ${\boldsymbol{\nabla}}$; and $\eta=\chi^{2}$ is the particle number density, as usual.
Consider a thin spherical shell defined by the radial interval from $\xi$ to $\xi+\delta\xi$ for small $\delta\xi$. Within this shell, we admit small perturbations $\delta\eta$ in the particle density and $\delta{\bf v}$ in the collective velocity, added to the fixed background values $\eta_{0}$ and ${\bf v}_{0}=0$ with the subscript “0”. Further, we use Eqs. (54)-(55), linearized in $\delta\eta$ and $\delta{\bf v}$.
Differentiating the continuity equation and substituting the Euler equation to eliminate $\delta{\bf v}$, we arrive at
00 0 $=$ $\partial_{t}\left[\partial_{t}\delta\eta+\eta_{0}{\boldsymbol{\nabla}}\delta{\bf v}\right]$ (56)
$=$ $\partial^{2}_{t}\delta\eta-\eta_{0}\Delta_{\xi}\left(2W-\frac{1}{\eta_{0}}\Delta_{\xi}\delta\eta\right).$
This equation suggests that the background value $\eta_{0}$ evolves more slowly over time and space than the perturbation $\delta\eta$. In the absence of gravitation, it is possible to allow the homogeneity of matter, when $\eta_{0}$ is constant.
Since the interaction $W$ comprises both the gravitational term, governed by the parameter $A$, and the self-interaction, we expand it as a series, retaining only the linear corrections in $\delta\eta$:
$W=W_{0}+A\Delta^{-1}_{\xi}\delta\eta+c^{2}_{s}(\chi_{0})\frac{\delta\eta}{\eta_{0}},$ (57)
by using the formula for the sound speed $c_{s}$, see (41).
Assuming that inside the thin shell the gravitational potential ${\cal V}_{0}$ can be replaced by the Newtonian $\sim 1/\xi$, and the self-interaction $j_{0}(\chi_{0})$ remains constant, we derive the equation:
$\partial_{t}^{2}\delta\eta-2c^{2}_{s}(\chi_{0})\Delta_{\xi}\delta\eta-2A\eta_{0}\delta\eta+\Delta^{2}_{\xi}\delta\eta=0.$ (58)
It is worth noting other arguments in Ki03 ; FHWM that exclude the influence of ${\cal V}_{0}$ on the evolution of excitations.
Setting
$\delta\eta=\chi_{0}(\delta\chi_{k}+\delta\overline{\chi}_{k}),$ (59)
where $\delta\overline{\chi}_{k}$ means the complex conjugate of $\delta\chi_{k}$, we define the spherically symmetric form of excitation as
$\delta\chi_{k}=\chi_{+k}f_{k}(\xi,t)+\chi_{-k}\overline{f}_{k}(\xi,t),$ (60)
$f_{k}(\xi,t)=\frac{1}{\xi}\exp{(\mathrm{i}k\xi-\mathrm{i}\omega_{k}t)},$ (61)
where $\chi_{\pm k}$ are real amplitudes of incident and outgoing spherical waves; $\omega_{k}$ is their frequency indexed by the wave number $k$.
Substituting, we obtain the dispersion relation:
$\omega^{(\pm)}_{k}=\pm\sqrt{k^{4}+2k^{2}c^{2}_{s}(\chi_{0})-2A\eta_{0}}.$ (62)
Its dimensional form looks like
$\frac{E^{(\pm)}_{k}}{\hbar}=\pm\sqrt{\frac{\hbar^{2}}{4m^{2}}q^{4}+q^{2}v^{2}_{s}-4\pi G\rho},$ (63)
where the wave vector squared $q^{2}=k^{2}/r_{0}^{2}$ and the mass density $\rho=\rho_{0}\eta_{0}$. The dimensional sound speed squared is
$v^{2}_{s}=\frac{\hbar^{2}}{2m^{2}r_{0}^{2}}c^{2}_{s}.$ (64)
Note that the obtained form of dispersion relation is well known H2019 ; Ch20 . And the models differ in the properties of the sound speed $v^{2}_{s}$.
Let us first analyze the case without accounting for gravitation, when $A=0$ and the dispersion relation is reduced to the form:
$\omega_{k}^{(\pm)}=\pm\sqrt{k^{4}-k^{2}\chi_{0}j_{1}(\chi_{0})}.$ (65)
We immediately see that at the turning points in Fig. 3, when $\chi_{0}=a_{n}$ from the sequence $\{a_{n}\}$ in (32), the free-particle spectrum, $\omega_{k}^{(\pm)}=\pm k^{2}$, is recovered because $c^{2}_{s}(a_{n})=0$.
Since $j_{1}(\chi_{0})$ remains positive for $\chi_{0}\in(a_{2n-1};a_{2n})$, $n\in\mathbb{N}$, the excitation modes with $k^{2}<\chi_{0}j_{1}(\chi_{0})$ become unstable due to the imaginary $\omega_{k}^{(\pm)}$. These instability conditions are consistent with those predicted thermodynamically, in particular, for the segments $AB$ and $CD$ in Fig. 3. On the other hand, the physically constructive phase-1 and phase-2 correspond to $\chi_{0}$ in the intervals $(a_{2};a_{3})$ and $(a_{4};a_{5})$, where $j_{1}(\chi_{0})<0$ and $c^{2}_{s}>0$. Thereby, Eq. (65) also confirms the sustainability of these states.
Therefore, stable structures of axionlike DM are feasible at $\chi\in(a_{2n};a_{2n+1})$, i.e. at $\chi>\chi_{c}$. In this range, a nonperturbative approach becomes necessary.
For a nonzero parameter $A\ll 1$, suppression of the gravitational Jeans instability in (62) necessitates the introduction of a lower bound $k_{J}$ on the wave numbers $k$. However, due to the substantially inhomogeneous DM distribution and the need for a nonperturbative approach, the limiting size of a stable DM cloud cannot be reliably determined based on $k_{J}$. Therefore, studies using time-dependent numerical analysis are needed to refine Jeans limit in size and mass. For this reason, we adopt the known limits for cold boson stars BGZ84 ; UTB02 and conservatively estimate the maximum stable mass as $M\sim 0.6\hbar c/(Gm)$.
III Examining the model in
the context of NGC 2366
III.1 Solving the field equations and
fitting the rotation curve
Due to the nonlinearity of Eq. (19), the initial conditions of the spatial evolution of the field $\chi$ cannot be given arbitrarily, but must be determined. Requiring the absence of cuspidity at the origin, i.e. $\chi(0)<\infty$, $\chi^{\prime}(0)=0$, and $\chi^{\prime\prime}(0)<0$ for given $A$ and $\nu$, the finite initial value $\chi(0)=z$ should satisfy the equation:
$Az^{2}+(\nu-j_{0}(z))(\nu-\cos{z})=0.$ (66)
It is derived by substituting the approximation $\chi(\xi)=\chi(0)+\chi^{\prime\prime}(0)\xi^{2}/2$ for $\xi\to 0$ into (19)-(20) and by finding $\chi^{\prime\prime}(0)$ GN23 . Note that the condition $\chi^{\prime\prime}(0)<0$ is equivalent to requiring $\nu>j_{0}(z)$ and $\nu<\cos{z}$ for positive $\nu$.
Figure 5: Black curves are the solutions of Eq. (66) for $A=4.5\times 10^{-3}$. The red cross marks the initial condition for the blue curve in Fig. 4.
The form of Eq. (66) indicates that its solution $z$ exists for small $A$. For this reason, we take $A$ to be of the order $10^{-3}$. Physically, this means that gravitational interaction is significantly weaker than the self-interaction between particles. However, we immediately encounter the problem of the solution ambiguity as we see in Fig. 5. Moreover, the number of “islands”, as in Fig. 5, increases with decreasing $A$. As we shall see, this ambiguity leads to solutions with different thermodynamic properties.
The absence of any solution $z$ for a given pair $(A,\nu)$ means that $\chi(\xi)=0$ everywhere. Except for large $A$, this happens for $\nu$ outside the range $[\nu_{\min};\nu_{\max}]$ for each “island”, where $\nu_{\max}(A)$ and $\nu_{\max}(A)$ are also found numerically from (66). For $\nu_{\min}<\nu<\nu_{\max}$, two branches of $\chi_{0}(\nu)$ can occur, which indicate the existence of two regimes and possibility of a first-order phase transition.
Before theoretically studying the effects in DM, we test the model ability to describe observables. It is natural to focus on the rotation curves (RCs) of galaxies Craciun2019 ; Kun2020 .
The RC of a DM-dominated (dwarf) galaxy can be derived in the parametric form:
$V(r)=\{(r(\xi),\,V(\xi))\,|\,r=r_{0}\xi,\,V=V_{0}v(\xi)\},$ (67)
$v(\xi)=\sqrt{\frac{n(\xi)}{\xi}},\quad n(\xi)=\int_{0}^{\xi}\chi^{2}(s)s^{2}\mathrm{d}s,$ (68)
where $r_{0}$ and $V_{0}$ are dimensional scales; $n(\xi)$ defines the number (or mass) of particles inside a sphere of radius $\xi$.
Having obtained the optimal parameters $A$, $V_{0}$, and $r_{0}$, we can extract the dark axion mass $m$, the mass density scale $\rho_{0}$ and the central density $\rho_{c}=\rho_{0}\chi^{2}(0)$, the axion decay constant $f_{\text{a}}$ as
$m=\sqrt{\frac{A}{2}}\frac{\hbar}{V_{0}r_{0}},\quad\rho_{0}=\frac{V_{0}^{2}}{4\pi Gr_{0}^{2}},\quad f_{\text{a}}=\frac{m_{\text{Pl}}V_{0}^{2}}{\sqrt{2\pi A}},$ (69)
where $m_{\text{Pl}}\equiv\sqrt{\hbar c/G}=2.176434\times 10^{-8}$ kg.
Figure 6: Rotation curve for NGC 2366. Black dots corresponds the data from Oh08 . The solid curve is obtained on the base of the blue solution in Fig. 4.
It turns out that the blue curve in Fig. 4 can be used to reproduce the adapted RC of irregular dwarf galaxy NGC 2366 in Fig. 6 for $r_{0}=1.18$ kpc and $V_{0}=6.58$ km/s. Although we have achieved a good fit, a realistic picture requires at least taking into account the small contribution of the baryon distribution, which is here absorbed by the effective chemical potential $\nu$.
The result of fitting the RC leads to the following values of DM characteristics: the dark axion mass $m=0.1171\times 10^{-22}~{}\text{eV}c^{-2}$, the decay constant $f_{\text{a}}=3.4978\times 10^{19}$ eV, and the central mass density $\rho_{c}=0.1671\times 10^{-20}~{}\text{kg}/\text{m}^{3}$. Although the values of $m$ and $\rho_{c}$ appear to be an order of magnitude smaller than expected, they are not surprising given the known results KMT .
Note that various methods for estimating dark‐boson masses yield a wide, often order‐of‐magnitude, spread in their values Zim25 . The anisotropy of the CMB requires $m>10^{-24}~{}\text{eV}c^{-2}$ Ev1 , and taking into account weak gravitational lensing gives $m>10^{-23}~{}\text{eV}c^{-2}$ Ev2 . The Lyman-$\alpha$ forest simulations raise the lower bound to $m>10^{-20}~{}\text{eV}c^{-2}$ Ev3 . Thus, each model determines the mass that best suits a particular regime.
The boson mass $m$ allows us to roughly estimate the limiting mass of gravitationally stable DM sphere, $M_{c}\sim 0.6\hbar c/(Gm)\simeq 6.84\times 10^{13}M_{\odot}$, while the computed DM halo mass is about $10^{10}M_{\odot}<M_{c}$, where $M_{\odot}$ is the solar mass. This confirms the gravitational stability of DM in NGC 2366.
Another fitting of the RC of NGC 2366 in the BEC DM concept is shown in Harko2011 . There, the positive $s$-scattering length $a_{s}$ and wave number $k$, $k^{2}=Gm^{3}/(\hbar^{2}a_{s})$, are used. In contrast, if the pairwise interaction $\psi^{4}$ is naively isolated from potential (3) in the vicinity of $\psi=0$, which corresponds to an unstable state, we obtain a negative scattering length from the relation:
$-\frac{Uv}{24}=\frac{2\pi a_{s}\hbar^{2}}{m}.$ (70)
Using it, we can formally write that
$A=\frac{Gm^{3}}{\hbar^{2}|a_{s}|}\frac{r_{0}^{2}}{6},$ (71)
to relate the dimensionless $A$ with $k$ in Harko2011 .
On the other hand, by expanding potential (3) in a series in the vicinity of some $\psi$ belonging to a stable state, for example, phase-1, one may obtain a positive scattering length $a_{s}>0$, as it should be.
Figure 7: Spatial dependence of the grand potential density (a) and the quantum potential (b) for the blue solution in Fig. 4.
It is useful to turn to the (shifted) grand canonical potential density $\omega$ and the potential $W$:
$\omega=(\partial_{\xi}\chi)^{2}-\nu\chi^{2}+A\Phi\chi^{2}-2\cos{\chi},$ (72)
$W(\xi)=A\Phi_{0}+A\Phi(\xi)+j_{0}(\chi(\xi)).$ (73)
The latter is involved in the Schrödinger-type equation $(-\Delta_{\xi}+W(\xi))\chi=\varepsilon\chi$, where $\varepsilon=1+2u$, and $\varepsilon\simeq-0.5867$ for the DM of NGC 2366.
Figure 8: Dependence of the grand potential density $\omega$ on order parameter $\chi$ for the blue solution in Fig. 4.
For the parameters and solution $\chi(\xi)$ leading to the RC of NGC 2366, the profiles of these potentials are shown in Fig. 7. Both indicate a potential well in the core region and have an unstable tail.
However, we wish to highlight a remarkable property of $\omega$ within the interval $6.32.5<\xi<$ 6.3, corresponding to the rarefied unstable phase. Specifically, within this band, the potential $\omega$ can take on the same value twice or even three times. This phenomenon is naturally linked to the repeated values of $\chi(\xi)$ in Fig. 4. The ambiguity is also visible in Fig. 8, where the curve segment $bc$ covers part of $ab$. It is natural to assume that, for certain $\chi$, the potential $\omega$ becomes ambiguous due to varying influences.
Thereby, we have identified transient processes occurring in the tail of the DM distribution for NGC 2366. Further, we explore these processes in more detail, focusing on solutions $\chi(\xi)$ similar to the blue one shown in Fig. 4, and using a statistical approach.
III.2 Analyzing the macroscopic state of BEC DM
Having extracted the model parameters from the galaxy’s rotation‐curve fit (e.g., NGC 2366), we then apply statistical analysis to probe the properties of BEC DM. Specifically, we focus on identifying the macroscopic state of DM in the galaxy. It depends on the fractions of the dense and rarefied phases of DM , and their change may be a phase transition. By mapping all configurations onto a diagram, the state of the galaxy would be represented by a single point there.
When analyzing the state of inhomogeneous matter, it is necessary to constrain its volume by imposing boundary conditions. The boundary radius $\xi_{B}$ of a spherical body in a vacuum is defined by the condition that the local pressure (43) vanishes, i.e. $p(\chi(\xi_{B}))=0$. Given that the solution $\chi(\xi)$ has no zeros for NGC 2366, we impose the condition $\chi(\xi_{B})=\pi$.
Figure 9: Spatial dependence of local pressure $p$ created by the blue solution $\chi(\xi)$ in Fig. 4 according to (75). The red dot marks the boundary of the system.
Based on the analysis of Figs. 4, 6 and 9, we choose $\xi_{B}\simeq 6.64$ to ensure $\chi(\xi_{B})=\pi$ and to cover almost the entire visible region of NGC 2366, which is bounded by $\xi\simeq 6.95$ in dimensionless units. At the same time, the presence of the cold DM CDM ; NFW beyond the galaxy is admitted, which is a common assumption in halo models.
When generating a set of solutions $\chi(\xi)$ for different values of $A$ and $\nu$ to support statistical analysis, we ensure that their shapes are similar to the reference solution shown in Fig. 4. To achieve this correspondence, varying $\nu$ at fixed $A$, we choose initial conditions from an “island” similar to that in Fig. 5 and impose certain conditions: $\chi^{\prime}(\xi_{B})>0$ and $\chi(\xi_{B})=\pi$, which should hold simultaneously and define $\xi_{B}$.
Each wave function $\chi(\xi)$ in this set defines a chemical potential $\mu$, an internal pressure $p$, and an internal energy density $e$ according to
$\mu=j_{0}(\chi),$ (74)
$p=\chi\sin{\chi}+2(1+\cos{\chi}),$ (75)
$e=-2(1+\cos{\chi}),$ (76)
which are connected by the Duhem–Gibbs and Euler relations (37). Thus, the $p-\mu$ diagram here is given by the diagram in Fig. 3 by shifting the origin.
On the other hand, Eq. (19) requires that the chemical potential $\mu$ be the sum of the fields:
$\mu=h-A\Phi,\quad h\equiv\nu+\frac{1}{\chi}\,\Delta_{\xi}\chi,$ (77)
where $h$ describes the long-wave quantum fluctuations LL .
The mean internal pressure induced by the axionlike self-interaction is defined as
$P=\frac{3}{\xi^{3}_{B}}\int_{0}^{\xi_{B}}p(\xi)\xi^{2}\,\mathrm{d}\xi,$ (78)
where $p(\xi)=p(\chi(\xi))$, substituting the Eq. (75).
Denoting the boundary value of the wave function as $\chi_{B}\equiv\chi(\xi_{B})$ and its derivative as $\chi^{\prime}_{B}\equiv\chi^{\prime}(\xi_{B})$, we proceed to analyze Eq. (78). Integrating by parts the trigonometric expressions, we arrive at
$P$ $=$ $-\chi_{B}\sin{\chi_{B}}+\frac{3}{\xi^{3}_{B}}\int_{0}^{\xi_{B}}\chi^{2}(\xi)\mu(\xi)\xi^{2}\,\mathrm{d}\xi$ (79)
$+\frac{2}{\xi^{3}_{B}}\int_{0}^{\xi_{B}}\chi^{\prime}(\xi)\chi(\xi)\mu(\xi)\xi^{3}\,\mathrm{d}\xi.$
The first term vanishes at $\chi_{B}=\pi$, but it may be needed in the general case.
Note that the functional (51) is then given by
$\frac{\Gamma}{\Gamma_{0}}=\frac{\xi_{B}^{3}}{3}\left(4-P+\frac{3}{\xi_{B}}\chi_{B}\chi^{\prime}_{B}\right).$ (80)
Using the Eq. (79), we can establish a relationship between the mean internal pressure and the partial pressures arising from quantum fluctuations and self-gravity by inserting the chemical potential from Eq. (77). Integrating, we obtain
$P$ $=$ $-\chi_{B}\sin{\chi_{B}}+\Pi_{\rm q}+A\Pi_{\rm gr},$ (81)
$\Pi_{\rm q}$ $=$ $\frac{3}{\xi_{B}}\chi_{B}\chi^{\prime}_{B}+(\chi^{\prime}_{B})^{2}+\nu\chi^{2}_{B}$ (82)
$-\frac{2}{\xi^{3}_{B}}\int_{0}^{\xi_{B}}(\chi^{\prime}(\xi))^{2}\xi^{2}\,\mathrm{d}\xi,$
$\Pi_{\rm gr}$ $=$ $-\Phi(\xi_{B})\chi^{2}_{B}+\frac{1}{\xi^{3}_{B}}\int_{0}^{\xi_{B}}\frac{n(\xi)\,\mathrm{d}n(\xi)}{\xi},$ (83)
where $n(\xi)$ is defined in (68). Since $\Pi_{\rm gr}<0$, one gets that $\Pi_{\rm q}>P>0$.
Thus, we extract the contribution $\Pi_{\rm q}$ of quantum fluctuations by setting $A=0$ in (81) for a moment and thereby dropping the self-gravity term $\Pi_{\rm gr}$. With increasing $\Pi_{\rm q}$, i.e. with enhancing fluctuations, it is expected an increase in the system size $\xi_{B}$ with a decrease in the particle number density
$\sigma=\frac{3}{\xi^{3}_{B}}N;\quad N=n(\xi_{B}),$ (84)
where $N$ defines the total number of particles.
At this stage, our goal is to determine under what conditions such characteristics continuously change and when the phase transition occurs. At zero temperature, this can be determined by varying the gravitational parameter $A$.
Figure 10: The BEC DM characteristics with changing pressure of quantum fluctuations $\Pi_{\rm q}$ at $A=4.5\times 10^{-3}$ (blue curves), $A=5.1\times 10^{-3}$ (red curves), $A=5.6\times 10^{-3}$ (green curves). The red dot corresponds to the solution in Fig. 4 applied to describe RC of NGC 2366 in Fig. 6.
The numerical results in Fig. 10 indicate that DM configurations with denser content continuously transform into a state with less dense content (and vice versa) when $A<5.1\times 10^{-3}$, while the first-order phase transition occurs at $A>5.1\times 10^{-3}$. This implies that the value of $A=5.1\times 10^{-3}$ is close to critical.
[…]
Although the basic macroscopic functions at $A=5.6\times 10^{-3}$, including the mean internal pressure $P$, are single-valued functions of $\nu$, the function $\Pi_{\rm q}$ is not. For example, the value $\Pi_{\rm q}=6.60$ occurs three times: at $\nu_{1}\simeq 0.493$, $\nu_{2}\simeq 0.611$, and $\nu_{3}\simeq 0.651$, see Fig. 10(a). Then the pressure jump in Fig. 10(e) during the phase transition is equal to the difference $\Delta P\equiv P(\nu_{3})-P(\nu_{1})=A(\Pi_{\rm gr}(\nu_{3})-\Pi_{\rm gr}(\nu_{1}))$ according to (81) for $\chi_{B}=\pi$. The positivity of $\Delta P\simeq 0.492$ indicates a weakening of the gravitational interaction. The jumps of other functions should then be expressed through $\Delta P$, which is precisely determined by applying Maxwell’s rule.
Note that this property of $\Pi_{\rm q}$ arises due to the long tail of the rarefied unstable phase, as in Fig. 4, which is defined by $\chi<\chi_{c}$.
Denoting $\chi_{0}\equiv\chi(0)$ and expanding the set of previously introduced characteristics, we included the mean energy density ${\cal E}$ in Fig. 10(f):
${\cal E}=\frac{3}{\xi^{3}_{B}}\int_{0}^{\xi_{B}}e(\xi)\xi^{2}\mathrm{d}\xi,$ (85)
and the DM phase-1 fraction in Fig. 10(h):
$x=\frac{100\%}{N}\int_{0}^{\xi_{B}}\theta(\chi(\xi)-\chi_{c})\chi^{2}(\xi)\xi^{2}\mathrm{d}\xi,$ (86)
where $\theta(z)$ is the Heaviside function.
Thus, an increase in quantum fluctuations $\Pi_{\rm q}$ predictably leads to an expansion of the system, see Fig. 10(g). Consequently, the mean densities of particle number and internal energy naturally decrease, as shown in Fig. 10(d) and (f). Additionally, the enhanced fluctuations inhibit the increase of the central particle density, which is proportional to $\chi^{2}_{0}$, see Fig. 10(c). Moreover, the transition to the predominance of the rarefied phase is accompanied by an increase in chemical potential $\nu$, see Fig. 10(a). However, the decreasing part of the chemical potential $u$ in Fig. 10(b) indicates the unstable state. The near-linear dependence on larger $\Pi_{\rm q}$ confirms this.
Physically, we observe that a reduction in the amount of phase-1, see Fig. 3, mainly concentrated in the core, results in a weakened gravitation, which, in turn, facilitates the growth of quantum fluctuations. On the contrary, the “massive” core ensures the stability of the entire system.
Note that there is confirmation of similar changes in properties and phase transition under the influence of “compression” $\Pi_{\nu}$, used in GKN20 :
$\Pi_{\nu}=-\frac{3}{\xi^{3}_{B}}\int_{0}^{\xi_{B}}\chi^{2}(\xi)h(\xi)\xi^{2}\mathrm{d}\xi,$ (87)
where $h(\xi)$ is given by (77). However, an increase in $\Pi_{\nu}$ corresponds to a decrease in $\Pi_{\rm q}$ and vice versa. We do not show these results here.
Figure 11: Total number of DM particles $N$ as a function of radius $\xi_{B}$ (a) and quantum fluctuation pressure $\Pi_{\rm q}$ (b) for $A=4.5\times 10^{-3}$ (blue curves), $A=5.1\times 10^{-3}$ (red curves), $A=5.6\times 10^{-3}$ (green curves). The red dot reflects the configuration of NGC 2366.
Valuable information can be also extracted from the mass-radius relation $M(R)$ by associating the total mass $M=4\pi r_{0}^{3}\rho_{0}N$ and the radius $R=r_{0}\xi_{B}$ with dimensionless quantities in Fig. 11(a). Although the number $N$ in the case of the unstable rarefied phase dominance tends to increase rapidly with radius $\xi_{B}>8$, we note the presence of a minimum $N_{\min}$ for relatively dense and compact objects at $\xi_{B}<8$. The existence of a minimum point separating the stable and unstable branches in our approach reproduces the predictions for the axiton BZ19 ; KT94 ; ax5 .
This minimum is also seen in Fig. 11(b) at $\Pi_{\rm q}\simeq 6.15$. Moreover, the total number (or mass) of particles can exhibit either continuous or discontinuous behavior during the first-order phase transition, depending on the value of $A$ in Fig. 11(b).
In Fig. 11(a) there are also intersections of curves with different $A$. It is noteworthy that the object configurations with different spatial distributions that share the same $N$ and $\xi_{B}$ correspond to the metastability regions in Fig. 11(b).
The BEC DM state of NGC 2366, marked by a red dot in Fig. 10 and 11 at $\Pi_{\rm q}\simeq 5.852$, is characterized by the presence of $x\simeq 19.2\%$ of the dense DM in Fig. 10(h), which is concentrated in $4.7\%$ of the volume. According to our model, it belongs to the (blue) continuous trajectory, which is classified as supercritical fluid with respect to the variations in $(A,\Pi_{\rm q})$. However, this conclusion relies on spatially averaged values and overlooks the inhomogeneous DM distribution, as evidenced in Fig. 4 and 9. In addition, the phase trajectory of a macroscopic system can be significantly altered by random or external influences. In fact, BEC DM of NGC 2366 exhibits a structured, fruitlike composition, consisting of a dense core, sparse pulp, and crust. These zones are defined by distinct radial intervals: 1) $r\leq 2.83$ kpc, including a central region for $r\leq 0.94$ kpc with a density ratio $\chi_{0}^{2}/\sigma\simeq 6.04$; 2) $2.83\leq r\leq 5.55$ kpc; 3) $5.55\leq r\leq 7.79$ kpc, representing the outer layer in the model.
Figure 12: Grand canonical potential $\Omega$ as a function of quantum fluctuation pressure $\Pi_{\rm q}$ for $A=4.5\times 10^{-3}$ (blue curve), $A=5.1\times 10^{-3}$ (red curve), $A=5.6\times 10^{-3}$ (green curve). The red dot corresponds to the parameters of NGC 2366.
In Fig. 12, the behavior of grand canonical potential, $\Omega=-(\xi_{B}^{3}/3)P$, shows that the DM of NGC 2366 is in a nearly stable state – its value of $\Omega$ remains on a plateau even with small fluctuations in its characteristics. This may explain the location of the galaxy to the left of the critical minimum $N_{\min}$ in Fig. 11.
Figure 13: Relations between gravitation properties and dense DM fraction $x$ for $A=4.5\times 10^{-3}$ (blue curves), $A=5.1\times 10^{-3}$ (red curves), $A=5.6\times 10^{-3}$ (green curves). Panel (a) illustrates the correlation between $x$ and the gravitational potential $N/\xi_{B}$. Panel (b) depicts the dependence of $x$ on dimensionless surface gravity ${\rm g}/A=N/\xi^{2}_{B}$.
On the other hand, a sharp decrease in $\Omega$ signals a fall into an unstable state.
The parameters that indicate the end of the stability plateau and the beginning of the instability regime can be associated with the minimum gravitational potential at the boundary sphere of radius $\xi_{B}$ in Fig. 13(a). The turning points between stable and unstable configurations, based on the analysis of three curves, are as follows: for the blue curve, $x\simeq 12.5\%$ and $\Pi_{\rm q}\simeq 6.32$; for the red curve, $x\simeq 12\%$ and $\Pi_{\rm q}\simeq 6.49$; and for the green curve, $x\simeq 11\%$ and $\Pi_{\rm q}\simeq 6.6$. In Fig. 13(a) the branches of the curves to the right of the minima show enhanced gravitational attraction due to a growing fraction of the dense DM component, while the branches to the left indicate intensified gravitation arising from the accumulation of DM in the unstable state as the segment $AB$ in Fig. 3.
In Fig. 13(b), it is shown that the impact of surface gravity, expressed as ${\rm g}=AN/\xi^{2}_{B}$, also confirms the presence of a first-order phase transition in the model, although ${\rm g}$ is not a thermodynamic function (potential) used to extract critical parameters. This confirms that its value becomes larger as $x$ increases.
IV Conclusion
We developed a statistical approach to the stationary nonrelativistic BEC DM model with instanton self-interaction, $V_{\text{inst}}=p_{0}-2\cos{\chi}$ in terms of the dimensionless real amplitude $\chi$ of the macroscopic wave function at $p_{0}=\pm 2$. Introducing the chemical potential $\mu$ and the internal pressure $p$ at zero temperature [see (35) and (43)], we classified stable, unstable and metastable states of matter using the $p-\mu$ phase diagram in the range $\chi\in[0;4.5\pi]$ in Fig. 3. The phase boundaries $\chi=a_{n}$ arise from the zeros of the spherical Bessel function, $j_{1}(a_{n})=0$ with $a_{1}=0$, defining intervals of stability $\chi\in(a_{2n};a_{2n+1})$ and instability $\chi\in(a_{2n-1};a_{2n})$ for $n\in\mathbb{N}$. Both the thermodynamic functions and the excitation spectrum confirm this partitioning, indicating that the squared sound speed $c^{2}_{s}(\chi)$ in (41) vanishes at $\chi=a_{n}$, is positive within stable bands, and negative within unstable ones. In particular, the small-amplitude regime at $\chi\in[0;\chi_{c}]$, where $\chi_{c}=a_{2}\simeq 4.493409$, is also shown to be unstable, which is consistent with earlier studies Nambo24 ; KT94 .
We argued that a first-order phase transition is possible between two existing stable phases with higher $\chi$ in the bands we have defined. Such amplitudes are usually ignored in the axion theory.
These findings explain why, taking into account self-gravity, the numerical solutions of the Gross–Pitaevskii–Poisson equations in the Thomas–Fermi approximation must be piecewise: radial integration necessarily stops at each “critical” amplitude $\chi=a_{n}$, see (48).
We apply the BEC model (19)-(20), which takes into account self-gravity and quantum fluctuations, to nonperturbatively describe DM in dwarf galaxies over the range $\chi\in[0;2.5\pi]$, which spans exactly one stable and one unstable phase. After prescribing mathematically consistent initial conditions for the nonlinear spatial evolution, we fit a representative numerical solution to the rotation curve of the DM-dominated dwarf NGC 2366 Oh08 . The fit yields a dark boson mass $m\simeq 0.1171\times 10^{-22}~{}\text{eV}c^{-2}$, which is comparable to the mass in KMT . While NGC 2366 exhibits active star formation and ISM reionization, we employ its DM configuration solely as a prototypical case of our statistical ensemble of DM distributions in similar galactic settings.
Requiring that the stable phase forms the dense solitonic core of the DM halo, we compute thermodynamic functions for NGC 2366 and a family of hypothetical analogues with different core-to-halo mass ratios. By varying the partial pressure induced by quantum fluctuations $\Pi_{\rm q}$, we trace the resulting macroscopic characteristics in Fig. 10, which undergo a first-order phase transition for certain strengths of the gravitational parameter $A$, the only coupling constant in the model. The revealed jumplike behavior of the functions is attributed to the ambiguity in the quantum pressure $\Pi_{\rm q}$.
In the case of NGC 2366, however, the DM lies in the supercritical regime, when its properties vary smoothly. Due to Fig. 10(h) and 13(a), we find that a core exceeding the threshold $\gtrsim 12\%$ of the total mass enhances gravity and suppresses quantum fluctuations in the surrounding rarefied halo, while the internal pressure $p$ ensures core stability. Concretely, NGC 2366’s DM dense core contains $\approx 19\%$ of the DM mass but occupies only $\approx 4.7\%$ of its volume.
We define the threshold of the core mass from the gravitational potential extremum in Fig. 13(a), comparing it with the end of the stability plateau of the grand canonical potential $\Omega$ in Fig. 12. The extremum existence means that the gravitational potential values are repeated for different configurations, thereby causing ambiguity in thermodynamic potential density $\omega$ in Fig. 8.
Moreover, in the mass-radius diagram in Fig. 11(a) – which shows stable and unstable branches analogous to those of the axiton BZ19 ; KT94 ; ax5 – the DM configuration for NGC 2366 lies firmly on the stable branch, consistent with its existence in nature.
Thus, the thermodynamic signature provides a novel diagnostic for distinguishing stable and unstable DM configurations in galactic halos. Viewed ergodically, it provides an alternative to time-dependent approaches, and a direct comparison between the two deepens physical insight.
Acknowledgements.
This work is supported by the Ukrainian-Swiss project No. IZURZ2-224868 “Compact star-forming galaxies and their impact on cosmic reionization and cosmology”.